OpenStudy (anonymous):

find the quotient of 20a^b+15ab^4-25ab/5ab

OpenStudy (luigi0210):

Is that a^b..?

OpenStudy (anonymous):

no

OpenStudy (luigi0210):

or is it 20ab?

OpenStudy (anonymous):

\[20a^3b+15ab^4-25ab/\]

OpenStudy (anonymous):

/5ab

OpenStudy (luigi0210):

\[\frac{ 20a^3b+15ab^4-25ab }{ 5ab }=\frac{ 20a^3b }{ 5ab }+\frac{ 15ab^4 }{ 5ab }-\frac{ 25ab }{ 5ab }\]

OpenStudy (anonymous):

so its 4a^2b+3b^3-5

OpenStudy (luigi0210):

Yup! :P